OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make.

OpenStudy (anonymous):

whats the function?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

f(x) = 15x + 85

OpenStudy (anonymous):

bracelet=$85 each charm=$15 85+X(15)=Y 85+(5)(15)= 85+75= 160 $160

OpenStudy (anonymous):

Awee thanks dear! so much!

OpenStudy (anonymous):

no problem